# Creation of domain-wall Skyrmions in chiral magnets

### v2

@@ 2 The chiral magnetic model
where we have rescaled length and energy units so as to only have dependence on one physical parameter, $\kappa$ [^1] The physical constants of ref. [57] correspond to the value of the coupling $\kappa\approx 0.4$. The field $n$ is interpreted as the magnetization vector in chiral magnets, but here we take it as a nonlinear sigma model field with target space $S^{2}$: $\mathbf{n}=(n_{1},n_{2},n_{3})$ satisfying $\mathbf{n}\cdot\mathbf{n}=1$.
@@ 2.1 Setup and initial state
where $E_{0}>0$. This is only possible for when the DMI energy is negative and hence when $\beta=\pi/2$ in our notation.
@@ 2.2.2 Phase diagram and final states
Fig. 3: The phase diagram of final states as a function of the initial parameters (a) $(X_{0},\alpha)$ for $\beta=\frac{\pi}{2}$ and (b) $(X_{0},\beta)$ for $\alpha=\frac{\pi}{2}$. The states are color coded with the same colors as in fig. 2: DW + Skyrmion (blue), DW-Skyrmion (green), only DW (red), anti-DW-Skyrmion and Skyrmion pair that annihilates and leaves behind two DW-Skyrmions (yellow) and finally the same as the previous case with the anti-DW-Skyrmion and bulk Skyrmion being a metastable bound state (magenta). The magenta state eventually decays into the yellow state and the yellow state can decay into the red state for finite specimen sizes. The marked points with (e.g. A1,B1$\cdots$) correspond to videos available in the supplemental material.
@@
$\ \ \mathbf{n}$Refer to caption Refer to caption Refer to caption $\mathcal{E}_{\rm DMI}$efer to caption Refer to caption Refer to caption
@@ 2.3 Interactions
The aim of this section is to understand, qualitatively, why the stable DW [$\alpha=\frac{\pi}{2}$] and the stable (isolated) Skyrmion [$\beta=\frac{\pi}{2}$] repel each other at large separation distances. This can be understood by studying the asymptotic interactions.
@@
$=\frac{1}{2}\partial_{i}\mathbf{n}\cdot\partial_{i}\mathbf{n}+\kappa\sum_{a=1}^{2}\epsilon^{aib}n^{a}\partial_{i}n^{b}-\kappa\epsilon^{ib3}\partial_{i}n^{3}n^{b}+\frac{1}{2}\left[1-(n^{3})^{2}\right]+\kappa\epsilon^{ib3}\partial_{i}(n^{3}n^{b}).$ (2.17)
@@
$\mathcal{E}$ $=\mathcal{E}_{0}-\frac{2\,\overline{\mathrm{eom}}}{(1+|u|^{2})^{3}}\delta u+\partial_{i}\left(\left[\frac{2\partial_{i}\bar{u}}{(1+|u|^{2})^{2}}-\frac{4\mathrm{i}\kappa(\bar{v}_{i}u-v_{i}\bar{u})}{(1+|u|^{2})^{3}}\right]\delta u\right)+\textrm{c.c.}+\mathcal{O}(\delta u^{2},\delta\bar{u}^{2}).$ (2.18)
The equation of motion (eom) vanishes on the background solution. Hence, to linear order we obtain [^2] Changing the phase $\alpha$ for the DW or the phase $\beta$ for the Skyrmion away from the stable point $(\alpha=\beta=\frac{\pi}{2})$ corresponds to turning on a quadratic field operator in the piece eom, which we neglect to leading order.
@@ 3 Conclusion and outlook
With our proposals, we have found two ways to capture an isolated Skyrmion into a DW, i.e. by rotating the DW phase or by rotating the Skyrmion on the target space, but there could be many more ways to perform a successful capture – some of all these could become experimentally viable in real chiral magnets. Although we performed the analysis using the Bloch DMI by convention, we could equivalently have used a Néel DM instead (see Appendix A).
@@
After posting this paper on the arXiv, a preprint with some overlap also appeared on the arXiv [65].
@@ Appendix A Bloch versus Néel DW-Skyrmion
+$\phantom{=\ }+\begin{cases}\kappa\mathbf{n}\cdot\nabla\times\mathbf{n},&\textrm{Dresselhaus SOC},\\ \kappa\left(\mathbf{n}\cdot\nabla n^{3}-n^{3}\nabla\cdot\mathbf{n}\right),\qquad&\textrm{Rashba SOC}.\end{cases}$ (A.5)
For the gauge field corresponding to the Dresselhaus SOC, the Lagrangian is simply given by eq. (A.1) up to a constant ($\kappa^{2}$) with $12m^{2}-\kappa^{2}=$ 1. The other gauge field gives rise to a seemingly different form of the DM term. It is known that the Dresselhaus-SOC-type DM term gives rise to the Bloch-type DW-Skyrmion, whereas the Rashba-SOC-type DM term gives rise to the Néel-type [46]. On the other hand, in the energy functional (A.1), the difference between the two types of DM term is simply a gauge choice. In fact, a gauge transformation

### v3

@@ 2 The chiral magnetic model
where we have rescaled length and energy units so as to only have dependence on one physical parameter, $\kappa$. [^1] The physical constants of ref. [57] correspond to the value of the coupling $\kappa\approx 0.4$. The field $n$ is interpreted as the magnetization vector in chiral magnets, but here we take it as a nonlinear sigma model field with target space $S^{2}$: $\mathbf{n}=(n_{1},n_{2},n_{3})$ satisfying $\mathbf{n}\cdot\mathbf{n}=1$.
@@ 2.1 Setup and initial state
where $E_{0}>0$. This is only possible when the DMI energy is negative and hence when $\beta=\pi/2$ in our notation.
@@ 2.2.2 Phase diagram and final states
[…]
@@
[…]
@@ 2.3 Interactions
The aim of this section is to understand, qualitatively, why the stable DW $\alpha=\frac{\pi}{2}$] and the stable (isolated) Skyrmion $\beta=\frac{\pi}{2}$] repel each other at large separation distances. This can be understood by studying the asymptotic interactions.
@@
$=\frac{1}{2}\partial_{i}\mathbf{n}\cdot\partial_{i}\mathbf{n}+2\kappa\epsilon^{ij}n^{i}\partial_{j}n^{3}+\frac{1}{2}\left[1-(n^{3})^{2}\right]+\kappa\epsilon^{ij}\partial_{i}(n^{j}n^{3}),\qquad i,j=1,2.$ (2.17)
@@
$\mathcal{E}$ $=\mathcal{E}_{0}-\frac{2\,\overline{\mathrm{eom}}}{(1+|u|^{2})^{3}}\delta u+\partial_{i}\left(\left[\frac{2\partial_{i}\bar{u}}{(1+|u|^{2})^{2}}-\frac{4\mathrm{i}\kappa(\bar{v}_{i}u-v_{i}\bar{u})\bar{u}}{(1+|u|^{2})^{3}}\right]\delta u\right)+\textrm{c.c.}+\mathcal{O}(\delta u^{2},\delta\bar{u}^{2}),$ (2.18)
where $v=(1,\mathrm{i})$. The equation of motion (eom) vanishes on the background solution. Hence, to linear order we obtain [^2] Changing the phase $\alpha$ for the DW or the phase $\beta$ for the Skyrmion away from the stable point $(\alpha=\beta=\frac{\pi}{2})$ corresponds to turning on a quadratic field operator in the piece eom, which we neglect to leading order.
@@ 3 Conclusion and outlook
With our proposals, we have found two ways to capture an isolated Skyrmion into a DW, i.e. by rotating the DW phase or by rotating the Skyrmion on the target space, but there could be many more ways to perform a successful capture – some of all these could become experimentally viable in real chiral magnets. Although we performed the analysis using the Bloch DMI by convention, we could equivalently have used a Néel DMI instead (see Appendix A).
@@
After posting this paper on the arXiv, a preprint with some overlap also appeared on the arXiv [65]. In Ref. [65], configurations similar to those we display in Fig. 2 are found and a configuration with two DWs is also studied. Instead, in our paper we provide a systematic result in form of the phase diagram shown in Fig. 3 and we compute the asymptotic interactions between the DW and the isolated magnetic Skyrmion analytically.
@@ Appendix A Bloch versus Néel DW-Skyrmion
$\phantom{=\ }+\begin{cases}\kappa\mathbf{n}\cdot\nabla\times\mathbf{n},&\textrm{Dresselhaus SOC},\\ \kappa\left(\mathbf{n}\cdot\nabla n^{3}-n^{3}\nabla\cdot\mathbf{n}\right),\qquad&\textrm{Rashba SOC}.\end{cases}$ (A.5)
For the gauge field corresponding to the Dresselhaus SOC, the Lagrangian is simply given by eq. (A.1) up to a constant ($\kappa^{2}$) with $2m^{2}-\kappa^{2}=1$. The other gauge field gives rise to a seemingly different form of the DM term. It is known that the Dresselhaus-SOC-type DM term gives rise to the Bloch-type DW-Skyrmion, whereas the Rashba-SOC-type DM term gives rise to the Néel-type [46]. On the other hand, in the energy functional (A.1), the difference between the two types of DM term is simply a gauge choice. In fact, a gauge transformation